OpenStudy (anonymous):

how would you solve √x2-x-6=x+3 (the xsquared and x and 6 are all under the radical

OpenStudy (anonymous):

squaring on both sides => x2 - x - 6 = x2 + 6x + 9 => 7x = -15 => x = -15/7

OpenStudy (anonymous):

the answer asks for a solution set so would i just write -15/7?

OpenStudy (precal):

\[\left\{x such that x=-.15/7 \right\}\] they are probably asking for set builder notation. There is a symbol for such that but I did not find it in the equation menu.

OpenStudy (anonymous):

it just asks what is the solution set and to separate my answer with a comma